In a particular country in 2008, the average worker had to work 40 hours to produce 100 units of output. In that same country in 2018, the average worker needed to work 35 hours to produce 94.5 units of output. In that country, the productivity of the average worker

a.

increased by 2 percent between 2008 and 2018.

b.

increased by 8 percent between 2008 and 2018.

c.

remained unchanged between 2008 and 2018.

d.

decreased by 2 percent between 2008 and 2018.
Answer #1

Solution: increased by 8 percent between 2008 and 2018

Explanation:

Hours

Units of output

Per unit output

2008

40

100

100/ 40 = 2.5

2018

35

94.5

94.5/35 = 2.7

Increase in output = (2.7 - 2.5) / 2.5 * 100 = 8%